# Enhancing Transformers for Generalizable First-Order Logical Entailment

### v1

@@ Appendix A Dataset Statistics
In this section, we introduce the statistics and query types of our benchmark dataset. Details of the query statistics for our dataset are presented in Table 9. For each knowledge graph, we sampled all one-hop projection (1p) queries from the training graph and sampled twice as many for the other 22 in-distribution query types. For validation and testing queries, we set the quantities to 8000, 5000, and 4000, respectively, following the convention established in KG literature Ren et al. (2020); Ren and Leskovec (2020); Bai et al. (2023b).
@@
Table 10 and 11 presented the details of query types in our benchmark. All 32 unseen query types are crafted compositionally based on the 23 seen query types. Table 12 presented the subset of 25 lisp-like compatible query types used in our experiment on query syntax.
@@
Appendix B Baseline Models
Appendix C Query Reversion
Type Name Before Reversion After Reversion
2p (r1(s1,e1))&(r2(e1,f)) (r2(e1,f))&(r1(s1,e1))
2in (r1(s1,f))&(!(r2(s2,f))) (!(r2(s2,f)))&(r1(s1,f))
Appendix D Attention Visualization
@@ Appendix G Evaluation Details
Appendix H Technical Details
@@
Computational Resource: All transformers are trained on four NVIDIA A100 GPUs for two days, with a batch size of 1024.

### v3

@@ Appendix A Dataset Statistics
In this section, we introduce the knowledge graph details and the statistics and query types of our benchmark dataset.
Our benchmark incorporates three widely used knowledge graphs—FB15k, FB15k-237, and NELL995—derived from large-scale KGs to evaluate reasoning capabilities. FB15k and FB15k-237 originate from Freebase, a comprehensive database of structured general knowledge covering domains such as entertainment, sports, and geography, with entities linked by diverse relations (e.g., "born-in," "part-of"). NELL995 is derived from the Never-Ending Language Learning (NELL) system, which extracts structured facts from web text, focusing on real-world entities and relations.
Details of the query statistics for our dataset are presented in Table 9. For each knowledge graph, we sampled all one-hop projection (1p) queries from the training graph and sampled twice as many for the other 22 in-distribution query types. For validation and testing queries, we set the quantities to 8000, 5000, and 4000, respectively, following the convention established in KG literature Ren et al. (2020); Ren and Leskovec (2020); Bai et al. (2023b).
@@
Table 11 and 12 presented the details of query types in our benchmark. All 32 unseen query types are crafted compositionally based on the 23 seen query types. Table 13 presented the subset of 25 lisp-like compatible query types used in our experiment on query syntax.
Appendix B Baseline Models
Appendix C Query Reversion
Type Name Before Reversion After Reversion
2p (r1(s1,e1))&(r2(e1,f)) (r2(e1,f))&(r1(s1,e1))
2in (r1(s1,f))&(!(r2(s2,f))) (!(r2(s2,f)))&(r1(s1,f))
Appendix D Attention Visualization
@@ Appendix G Evaluation Details
Appendix H Transformer Architecture Comparison
We investigate the impact of different transformer architectures on complex query answering performance. Table 15 compares three architectural variants: encoder-only (our main approach), decoder-only, and encoder-decoder configurations. The results demonstrate that the encoder-only architecture with relative positional encoding achieves the best overall performance across datasets.
Dataset Architecture PE Type ID (Q) OOD (Q) All Queries
ID (K) OOD (K) ID (K) OOD (K) ID (K) OOD (K)
[…]
Encoder-only Relative PE 48.1 32.3 35.4 21.5 40.7 26.0
[…]
FB15k-237 Encoder-only Absolute PE 54.4 20.0 20.8 9.0 34.9 13.6
[…]
Decoder-only Absolute PE 55.2 19.9 21.9 9.8 35.8 14.0
[…]
Decoder-only Absolute PE 65.6 18.6 28.0 8.9 43.7 13.0
[…]
Table 15: Comparison of different transformer architectures (in MRR%).
Appendix I Technical Details
@@
Computational Resource: All transformers are trained on four NVIDIA A100 GPUs for two days, with a batch size of 1024. As shown in Table 16, TEGA achieves computational efficiency comparable to that of the baseline models.
@@
Models Inference Time (ms)
Transformer (Absolute PE) 43.2
Transformer (Disentangled PE) 47.3
Transformer (Relative PE) 44.7
+Adjacency Matrices Masking 46.8
+Direct Distance Encoding 44.9
TEGA (ours) 52.1
Table 16: Computational efficiency of TEGA compared with baselines.